# Hidden ferromagnetism of centrosymmetric antiferromagnets

### v1

@@ VI Examples
VI.1 Square lattice of perovskites
@@
As the first example, let us consider the single layer of orthorhombically distorted perovskites with the $Pbnm$ or $Bmab$ symmetry [26]. The lattice is specified by the primitive translations ${\bf T}_{1}=(1,0,0)$ and ${\bf T}_{2}=(0,1,0)$, and there are two sublattices, which are connected by the vectors ${\bf t}_{1}=(\frac{1}{2},\frac{1}{2})$ and ${\bf t}_{2}=(-\frac{1}{2},\frac{1}{2})$. The symmetry operation transforming two sublattices to each other is $\{{\cal C}_{2x}|{\bf t}_{1}\}$, meaning that $x$ components of the SO coupling is sign-alternating. For the space group $Bmab$, the lattice is additionally invariant under the mirror reflection $y\to-y$.
@@
The minimal model includes the following ingredients [26]: (i) The hoppings between first, second, and third nearest neighbors ($t_{1}$, $t_{2}$, and $t_{3}$, respectively); (ii) The orthorhombic strain of the second-nearest neighbor hoppings, $\delta t_{2}$, such that the total hopping along $x$ and $y$ is, respectively, $t_{2}+\delta t_{2}$ and $t_{2}-\delta t_{2}$ in both sublattices; (iii) The SO coupling in the noncentrosymmetric nearest-neighbor bonds, which can be taken in the form $\hat{\cal H}^{\rm so}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\pm it_{x}\hat{\sigma}_{x}$. The $y$ and $z$ components of $\hat{\cal H}^{\rm so}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, responsible for the weak spin ferromagnetism, can be eliminated via the unitary transformation [29, 50], as explained in Appendix A. We would like to reiterate again that the SO interaction obeys the following property, being the consequence of the inversion symmetry and translational invariance: $\hat{\cal H}^{\rm so}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}=(-1)^{m^{\prime\prime}+n^{\prime\prime}}\,\hat{\cal H}^{\rm so}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$. The regular SO interaction is not periodic on the lattice, specified by ${\bf t}_{1}$ and ${\bf t}_{2}$. However, it can be made periodic in the local coordinate frame as explained in Secs. II-IV; (iv) The Néel field, which is also parallel to $x$, $(-1)^{m+n+1}\,B\hat{\sigma}_{x}$, where odd and even values of mm++$n$ correspond to the sublattices $1$ and $2$, respectively. This term is transformed to $B\hat{\sigma}_{z}$ in the local coordinate frame, which is specified by the transformation (2) with $\alpha=0$ and $\boldsymbol{q}=(2\pi,0)$. We do not consider the altermagnetic deformation of the third-neighbor hoppings, which is identically equal to zero for the $Bmab$ symmetry of La2CuO4 and expected to be small for other materials [26].
@@
Thus, one can readily find the Hamiltonian (16) in the local coordinate frame, where $h^{\phantom{0}}_{\boldsymbol{k}}$, $h^{1}_{\boldsymbol{k}}$, and $h^{y}_{\boldsymbol{k}}$ are obtained by combining the Fourier transforms with the matrix elements of $\hat{\mathbb{1}}$ and $\hat{\sigma}_{x}$ given by Eqs. (10) and (7), respectively. Moreover, it is convenient to shift the $\boldsymbol{k}$-mesh: $(k_{x},k_{y})\to(k_{x}+\pi,k_{y})$. Altogether, this yields: $h^{\phantom{0}}_{\boldsymbol{k}}=2t_{2}(\cos k_{x}+\cos k_{y})+2\delta t_{2}(\cos k_{x}-\cos k_{y})+4t_{3}\cos k_{x}\cos k_{y}$, $h^{1}_{\boldsymbol{k}}=4t_{1}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$, and $h^{y}_{\boldsymbol{k}}=-4t_{x}\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}$. One can clearly see that, after replacing the Pauli matrices $\boldsymbol{\sigma}$ by pseudospin matrices $\boldsymbol{\tau}$, the obtained Hamiltonian is the same as the one in global coordinate frame for the $\sigma=$- states and $\delta t_{3}=0$, which is given by Eq. (1) of Ref. [26]. Obviously, it will yield the same anomalous Hall conductivity and orbital magnetization. Nevertheless, the Berry curvature can be now written in the more compact form, using Eq. (28):
@@
In comparison with calculations in the global coordinate frame [26], the generalized Bloch theorem allows us to use smaller unit cell. Therefore, the first Brillouin zone becomes larger, where $\Omega^{z}_{\boldsymbol{k}}$ has two positive and two negative areas [Fig. 4(a)], which tend to compensate each other. Therefore, in order to obtain finite $\sigma_{xy}$ and ${\cal M}^{z}$, it is essential to consider the orthorhombic strain, $\delta t_{2}$ [17, 26]. This strain deforms the Fermi surface, as shown in Fig. 4(e), so that the positive and negative areas of $\Omega^{z}_{\boldsymbol{k}}$ will contribute to $\sigma_{xy}$ with different weights, resulting in finite $\sigma_{xy}$. According to Eq. (38), in the case of ${\cal M}^{z}$, there is an additional modulation of $\Omega^{z}_{\boldsymbol{k}}$ by $h^{\phantom{0}}_{\boldsymbol{k}}$, which makes inequivalent the positive and negative areas of ${\cal M}^{z}_{\boldsymbol{k}}$ [see Fig. 4(b)]. Nevertheless is small if $\delta t_{2}$ is small.
@@ VI.2 VF4 and CuF2
The atoms in the nearest neighborhood of each atomic site are located at ${\bf t}_{3}$, -${\bf t}_{3}$, ${\bf t}_{3}$-${\bf T}_{2}$, and -${\bf t}_{3}$${\bf T}_{2}$. The AFM propagation vector can be taken as $\boldsymbol{q}=\boldsymbol{G}_{2}$, so that $\sin\frac{\boldsymbol{q}\cdot{\bf t}_{3}}{2}=1$ and $\sin\frac{\boldsymbol{q}\cdot({\bf t}_{3}-{\bf T}_{2})}{2}=-1$. Furthermore, in the first two bonds, the SO coupling is $it_{y}$, while in the second two bonds it is $it_{y}$. Then, using Eqs. (10) and (8) for $\alpha=\frac{\pi}{2}$, and shifting the $\boldsymbol{k}$-mesh $\boldsymbol{k}\to\boldsymbol{k}+\frac{\boldsymbol{q}}{2}$, one can find $h^{1}_{\boldsymbol{k}}=4t_{1}\cos\frac{\boldsymbol{k}\cdot({\bf T}_{1}-{\bf T}_{3})}{2}\cos\frac{\boldsymbol{k}\cdot{\bf T}_{2}}{2}$ and $h^{y}_{\boldsymbol{k}}=4t_{y}\sin\frac{\boldsymbol{k}\cdot({\bf T}_{1}-{\bf T}_{3})}{2}\sin\frac{\boldsymbol{k}\cdot{\bf T}_{2}}{2}$. Finally, $h^{\phantom{0}}_{\boldsymbol{k}}$ in Eq. (16) is the Fourier image of transfer integrals between atoms of one magnetic sublattice and assuming that the they are the same in both sublattices (otherwise the generalized Bloch theorem does not apply). Technically, $h^{\phantom{0}}_{\boldsymbol{k}}$ is obtained by averaging the transfer integrals over two magnetic sublattices. The details and applicability of this approximation and discussed in Appendix B.
@@
These tendencies are clearly seen in the behavior of AHE and orbital magnetization (see Fig. 5).
@@ VI.3 $\rm RuO_{2}$}-type systems
RuO2 is crystallized in body-centered tetragonal (bct) structure with two sublattices (the space group $P4_{2}/mnm$, No. 136). The atomic positions in both sublattices can be formally generated by the vectors ${\bf t}_{1}=(-\frac{1}{2},\frac{1}{2},\frac{1}{2})$, ${\bf t}_{2}=(\frac{1}{2},-\frac{1}{2},\frac{1}{2})$, and ${\bf t}_{3}=(\frac{1}{2},\frac{1}{2},-\frac{1}{2})$, in units of $a$ and $c$ ($c/a<1$ being the tetragonal distortion along $z$). The additional complication arising in this case is that there are two sign-alternating components of the SO coupling in the nearest bonds: $x$ and $y$, as explained in Fig. 6.
@@
The third ($z$) component is identically equal to zero. Thus, formally speaking, there is no weak spin ferromagnetism for the $P4_{2}/mnm$ symmetry.
@@
where it is again convenient to shift the $\boldsymbol{k}$-mesh: $\boldsymbol{k}\to\boldsymbol{k}+(\pi,0,0)$. Then, the parameters of Eq. (43) will be given by $h^{\phantom{0}}_{\boldsymbol{k}}=2(t_{2}-\delta t_{2})(\cos k_{x}+\cos k_{y})+2(t_{2}+\delta t_{2})\cos k_{z}$, $h^{1}_{\boldsymbol{k}}=8t_{1}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}$, $h^{y}_{\boldsymbol{k}}=-8t_{x}\sin\frac{k_{x}}{2}\cos\frac{k_{y}}{2}\sin\frac{k_{z}}{2}$, and $h^{z}_{\boldsymbol{k}}=8t_{y}\cos\frac{k_{x}}{2}\sin\frac{k_{y}}{2}\cos\frac{k_{z}}{2}$. Although for the $P4_{2}/mnm$ symmetry $t_{x}=t_{y}$, it is convenient to treat $t_{x}$ and $t_{y}$ as independent parameters. The eigenvalues, eigenfunctions, and properties of Hamiltonian (43) can be obtained along the same line as in Sec. V after replacing $B$ by $B+h^{z}_{\boldsymbol{k}}$.
Furthermore, we have the following properties: $h^{\phantom{0}}_{-\boldsymbol{k}}=h^{\phantom{0}}_{\boldsymbol{k}}$, $h^{1}_{-\boldsymbol{k}}=h^{1}_{\boldsymbol{k}}$, $h^{y}_{-\boldsymbol{k}}=h^{y}_{\boldsymbol{k}}$, and $h^{z}_{-\boldsymbol{k}}=-h^{z}_{\boldsymbol{k}}$, which immediately yields: $\varepsilon^{\pm}_{\boldsymbol{k}}(B)=\varepsilon^{\pm}_{-\boldsymbol{k}}(-B)$ and $\varepsilon^{\pm}_{\boldsymbol{k}}(B,t_{y})=\varepsilon^{\pm}_{\boldsymbol{k}}(-B,-t_{y})$. Then, the Berry curvature satisfies the following properties: $\Omega_{\boldsymbol{k}}(B)=-\Omega_{-\boldsymbol{k}}(-B)$ ($\partial_{k_{a}}\phi_{\boldsymbol{k}}$ changes the sign for $a=$ $z$ and $x$, while $\sin 2\theta_{\boldsymbol{k}}$ and $\partial_{k_{a}}\theta_{\boldsymbol{k}}$ do not change the sign) and $\Omega_{\boldsymbol{k}}(B,t_{y})=-\Omega_{\boldsymbol{k}}(-B,-t_{y})$ ($\partial_{k_{a}}\theta_{\boldsymbol{k}}$ changes the sign, while $\sin 2\theta_{\boldsymbol{k}}$ and $\partial_{k_{a}}\phi_{\boldsymbol{k}}$ do not change the sign). Therefore, the Hall conductivity, $\sigma_{zx}$, will be odd in $B$, $\sigma_{zx}(-B)=-\sigma_{zx}(B)$, and even in $t_{y}$, $\sigma_{zx}(-t_{y})=\sigma_{zx}(t_{y})$. Moreover, $\sigma_{zx}$ is odd in $t_{x}$: $\sigma_{zx}(-t_{x})=-\sigma_{zx}(t_{x})$ (i.e., for the SO component, which is parallel to the Néel field) [26]. Thus, in order to calculate $\sigma_{zx}$ to the first order in the SO coupling, $t_{y}$ can be neglected. This is consistent with the general strategy for calculating DM interactions using the response theory: in order to calculate some particular (say, $x$) component of the DM vectors, it is sufficient to align the exchange field along $x$ and consider only the $x$ component of the SO coupling [30, 59].

### v2

@@ VI Examples
VI.1 Square perovskite lattice
@@
As the first example, let us consider the single layer of orthorhombically distorted perovskites with the $Pbnm$ or $Bmab$ symmetry [26]. The lattice is specified by the primitive translations ${\bf T}_{1}=(1,0,0)$ and ${\bf T}_{2}=(0,1,0)$, and there are two sublattices, which are connected by the vectors ${\bf t}_{1}=(\frac{1}{2},\frac{1}{2})\equiv{\bf t}$ and ${\bf t}_{2}=(-\frac{1}{2},\frac{1}{2})$. The symmetry operation transforming two sublattices to each other is $\{{\cal C}_{2x}|{\bf t}\}$, meaning that $x$ components of the SO coupling is sign-alternating. For the space group $Bmab$, the lattice is additionally invariant under the mirror reflection $y\to-y$.
@@
The minimal model was introduced in Sec. V.1.1 and explained in Fig. 4. We would like to emphasize again that the SO interaction obeys the following property, being the consequence of the inversion symmetry and translational invariance: $\hat{\cal H}^{\rm so}_{\boldsymbol{R}+\boldsymbol{R}^{\prime\prime},\boldsymbol{R}^{\prime}+\boldsymbol{R}^{\prime\prime}}=(-1)^{m^{\prime\prime}+n^{\prime\prime}}\,\hat{\cal H}^{\rm so}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$. In the global frame the SO interaction is not periodic on the lattice, specified by ${\bf t}_{1}$ and ${\bf t}_{2}$. However, it can be made periodic in the local frame as explained in Secs. II-IV. We do not consider the altermagnetic deformation of the third-neighbor hoppings, which is identically equal to zero for the $Bmab$ symmetry of La2CuO4 and expected to be small for other materials [26].
@@
Thus, one can readily find the Hamiltonian (17) in the local coordinate frame, where $h^{\phantom{0}}_{\boldsymbol{k}}$, $h^{1}_{\boldsymbol{k}}$, and $h^{y}_{\boldsymbol{k}}$ are obtained by combining the Fourier transforms with the matrix elements of $\hat{\mathbb{1}}$ and $\hat{\sigma}_{x}$ given by Eqs. (10) and (7), respectively. Moreover, it is convenient to shift the $\boldsymbol{k}$-mesh: $(k_{x},k_{y})\to(k_{x}+\pi,k_{y})$. Altogether, this yields: $h^{\phantom{0}}_{\boldsymbol{k}}=2t_{2}(\cos k_{x}+\cos k_{y})+2\delta t_{2}(\cos k_{x}-\cos k_{y})+4t_{3}\cos k_{x}\cos k_{y}$, $h^{1}_{\boldsymbol{k}}=4t_{1}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$, and $h^{y}_{\boldsymbol{k}}=-4t_{x}\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}$, which coincide with the parameters reported in Ref. [26] in the global coordinate frame [53].
Berry curvature can be written in the compact form, using Eq. (29):
@@
In comparison with calculations in the global coordinate frame [26], the generalized Bloch theorem allows us to use smaller unit cell. Therefore, the first Brillouin zone becomes larger, where $\Omega^{z}_{\boldsymbol{k}}$ has two positive and two negative areas [Fig. 5(a)], which tend to compensate each other. Therefore, in order to obtain finite $\sigma_{xy}$ and ${\cal M}^{z}$, it is essential to consider the orthorhombic strain, $\delta t_{2}$ [17, 26]. Although the Berry curvature does not depend on $\delta t_{2}$, the latter deforms the Fermi surface, as shown in Fig. 5(e), so that the positive and negative areas of $\Omega^{z}_{\boldsymbol{k}}$ will contribute to $\sigma_{xy}$ with different weights, resulting in finite $\sigma_{xy}$. According to Eq. (39), in the case of ${\cal M}^{z}$, there is an additional modulation of $\Omega^{z}_{\boldsymbol{k}}$ by $h^{\phantom{0}}_{\boldsymbol{k}}$, which makes inequivalent the positive and negative areas of ${\cal M}^{z}_{\boldsymbol{k}}$ [see Fig. 5(b)]. Nevertheless the effect is small if $\delta t_{2}$ is small.
@@ VI.2 VF4 and CuF2
The atoms in the nearest neighborhood of each atomic site are located at ${\bf t}_{3}$, -${\bf t}_{3}$, ${\bf t}_{3}$-${\bf T}_{2}$, and -${\bf t}_{3}$${\bf T}_{2}$. The AFM propagation vector can be taken as $\boldsymbol{q}=\boldsymbol{G}_{2}$, so that $\sin\frac{\boldsymbol{q}\cdot{\bf t}_{3}}{2}=1$ and $\sin\frac{\boldsymbol{q}\cdot({\bf t}_{3}-{\bf T}_{2})}{2}=-1$. Furthermore, in the first two bonds, the SO coupling is $it_{y}$, while in the second two bonds it is $it_{y}$. Then, using Eqs. (10) and (8) for $\alpha=\pi/2$, and shifting the $\boldsymbol{k}$-mesh $\boldsymbol{k}\to\boldsymbol{k}+\frac{\boldsymbol{q}}{2}$, one can find $h^{1}_{\boldsymbol{k}}=4t_{1}\cos\frac{\boldsymbol{k}\cdot({\bf T}_{1}-{\bf T}_{3})}{2}\cos\frac{\boldsymbol{k}\cdot{\bf T}_{2}}{2}$ and $h^{y}_{\boldsymbol{k}}=4t_{y}\sin\frac{\boldsymbol{k}\cdot({\bf T}_{1}-{\bf T}_{3})}{2}\sin\frac{\boldsymbol{k}\cdot{\bf T}_{2}}{2}$. Finally, $h^{\phantom{0}}_{\boldsymbol{k}}$ in Eq. (17) is the Fourier image of transfer integrals between atoms of one magnetic sublattice and assuming that the they are the same in both sublattices (otherwise the generalized Bloch theorem does not apply). Technically, $h^{\phantom{0}}_{\boldsymbol{k}}$ is obtained by averaging the transfer integrals over two magnetic sublattices. The details and applicability of this approximation and discussed in Appendix B. Briefly, the hopping parameters $\delta t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ responsible for the altermagnetic spin-splitting of bands are small in VF4, but can be comparable with regular hoppings $t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ in the case of CuF2. In the strong coupling limit, the contribution of degenerate bands to AHE, stemming from the Berry curvature, is of the order of $\frac{t_{1}t_{y}}{B|B|}$, where $B\approx\frac{U}{2}$. The spin splitting yields the additional contribution $\sim\frac{2t_{1}t_{y}\delta t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}}{|B|^{3}}$, which is smaller by the factor $\frac{2\delta t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}}{|B|}$  [26]. Using realistic parameters $|\delta t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}|<30$ meV (Appendix B) and $U=4$ eV (Table 2), one can obtain the following estimate $\frac{2|\delta t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}|}{|B|}<0.03$. Therefore, the altermagnetic contribution to AHE, originating from the spin-splitting of bands, is expected to be small.
@@
These tendencies are clearly seen in the behavior of AHE and orbital magnetization (Fig. 6).
@@ VI.3 $\rm RuO_{2}$}-type systems
RuO2 is viewed as a canonical altermagnetic material [16]. It crystallizes in the rutile-type body-centered tetragonal (bct) structure with two sublattices (the space group $P4_{2}/mnm$, No. 136). The atomic positions in both sublattices can be formally generated by the vectors ${\bf t}_{1}=(-\frac{1}{2},\frac{1}{2},\frac{1}{2})$, ${\bf t}_{2}=(\frac{1}{2},-\frac{1}{2},\frac{1}{2})$, and ${\bf t}_{3}=(\frac{1}{2},\frac{1}{2},-\frac{1}{2})$, in units of $a$ and $c$ ($c/a<1$ being the tetragonal distortion along $z$). The additional complication arising in this case is that there are two sign-alternating components of the SO coupling in the nearest bonds: $x$ and $y$, as explained in Fig. 7.
@@
The third ($z$) component is identically equal to zero. Thus, formally speaking, there is no weak spin ferromagnetism arising from DM interactions, which is believed to be a unique aspect of RuO2 and other AFM materials with the $P4_{2}/mnm$ symmetry. However, in other materials, the weakly FM component can be eliminated in the local coordinate frame (see Appendix A). Then, from this point of view, there is no fundamental difference between RuO2 and other weak ferromagnets.
We would also like to note that the weak ferromagnetism in rutile compounds is allowed by symmetry [62] and the microscopic mechanism behind it is believed to be the single-ion anisotropy [63]. However, the single-ion anisotropy is even and at least of the second order in the SO coupling. Therefore, we expect that it will not contribute to AHE and orbital magnetization, which are odd in the SO coupling.
@@
where it is again convenient to shift the $\boldsymbol{k}$-mesh: $\boldsymbol{k}\to\boldsymbol{k}+(\pi,0,0)$. Then, the parameters of Eq. (44) will be given by $h^{\phantom{0}}_{\boldsymbol{k}}=2(t_{2}-\delta t_{2})(\cos k_{x}+\cos k_{y})+2(t_{2}+\delta t_{2})\cos k_{z}$, $h^{1}_{\boldsymbol{k}}=8t_{1}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}$, $h^{y}_{\boldsymbol{k}}=-8t_{x}\sin\frac{k_{x}}{2}\cos\frac{k_{y}}{2}\sin\frac{k_{z}}{2}$, and $h^{z}_{\boldsymbol{k}}=8t_{y}\cos\frac{k_{x}}{2}\sin\frac{k_{y}}{2}\cos\frac{k_{z}}{2}$, where $t_{1}$ is the nearest-neighbor hopping between sublattices and $t_{2}\pm\delta t_{2}$ are the hoppings within sublattices: along $z$ ($t_{2}+\delta t_{2}$) and in the $xy$ plane ($t_{2}-\delta t_{2}$). Although for the $P4_{2}/mnm$ symmetry $t_{x}=t_{y}$, it is convenient to treat $t_{x}$ and $t_{y}$ as independent SO parameters. The eigenvalues, eigenfunctions, and properties of Hamiltonian (44) can be obtained along the same line as in Sec. V after replacing $B$ by $B+h^{z}_{\boldsymbol{k}}$.
Furthermore, we have the following properties: $h^{\phantom{0}}_{-\boldsymbol{k}}=h^{\phantom{0}}_{\boldsymbol{k}}$, $h^{1}_{-\boldsymbol{k}}=h^{1}_{\boldsymbol{k}}$, $h^{y}_{-\boldsymbol{k}}=h^{y}_{\boldsymbol{k}}$, and $h^{z}_{-\boldsymbol{k}}=-h^{z}_{\boldsymbol{k}}$, which immediately yields: $\varepsilon^{\pm}_{\boldsymbol{k}}(B)=\varepsilon^{\pm}_{-\boldsymbol{k}}(-B)$ and $\varepsilon^{\pm}_{\boldsymbol{k}}(B,t_{y})=\varepsilon^{\pm}_{\boldsymbol{k}}(-B,-t_{y})$. Then, the Berry curvature satisfies the following properties: $\Omega^{y}_{\boldsymbol{k}}(B)=-\Omega^{y}_{-\boldsymbol{k}}(-B)$ ($\partial_{k_{a}}\phi_{\boldsymbol{k}}$ changes the sign for $a=$ $z$ and $x$, while $\sin 2\theta_{\boldsymbol{k}}$ and $\partial_{k_{a}}\theta_{\boldsymbol{k}}$ do not change the sign) and $\Omega^{y}_{\boldsymbol{k}}(B,t_{y})=-\Omega^{y}_{\boldsymbol{k}}(-B,-t_{y})$ ($\partial_{k_{a}}\theta_{\boldsymbol{k}}$ changes the sign, while $\sin 2\theta_{\boldsymbol{k}}$ and $\partial_{k_{a}}\phi_{\boldsymbol{k}}$ do not change the sign). Therefore, the Hall conductivity, $\sigma_{zx}$, will be odd in $B$, $\sigma_{zx}(-B)=-\sigma_{zx}(B)$, and even in $t_{y}$, $\sigma_{zx}(-t_{y})=\sigma_{zx}(t_{y})$. Moreover, $\sigma_{zx}$ is odd in $t_{x}$, i.e., when the SO coupling is in the direction of the Néel field: $\sigma_{zx}(-t_{x})=-\sigma_{zx}(t_{x})$. Thus, in order to calculate $\sigma_{zx}$ to the first order in the SO coupling, $t_{y}$ can be neglected. This is consistent with the general strategy for calculating DM interactions using the response theory: in order to calculate some particular (say, $x$) component of the DM vectors, it is sufficient to align the exchange field along $x$ and consider only the $x$ component of the SO coupling [30, 64].